# TDE-3: An improved prior for optical flow computation in spiking neural networks

### v1

@@ 5.2.1 Training and inference with the synthetic data
The SNNs were simulated using the PyTorch framework, inspired by previous implementations [37, 4]. Depending on the application, networks included several types of elements such as TDEs, spatiotemporal correlation filters (STCF), and low-pass filters of the spike trains (i.e. spike traces). Each element was defined by its synaptic weights and time constants. The TDE trainable parameters included gain, current, and voltage time constants and synaptic weight of the gain compartment. To enhance learning of the neuronal time constants, we used the approach developed by [47] that reformulated neuronal decays as the result of multiplying gain, current, and voltage with the Sigmoid of a trainable parameter.
InputptSpatio-TemporalCorrelation Filter TDEsVelocity = $\Sigma$ spikes
@@
Depending on the coding scheme (spike count or ISI), we used two slightly different training pipelines to extract motion direction and velocity (Figures 9,10). The two first steps are common to both pipelines. The input layer simply converts events to spikes. Then, these spikes are fed into the layer of TDE-3s, which effectively perform 1D convolution with kernel size=3, stride=1, and weights being facilitator, trigger, and inhibitor. This layer consists of two channels, one for each motion direction. At the next processing stage, spike count-based and ISI-based pipelines deviate. For the spike count-based velocity estimator we can straightaway find the total number of spikes emitted by each of the detectors, convert it to estimated velocity, calculate the loss, and perform a backward pass. For the ISI-based velocity estimation, the training of the network is not so straightforward and is described in section 5.2.2.
@@
In the experiments with noise (Figure 6, Table 2), not every increase in current is caused by a moving edge, and one of the goals of the training is to suppress the TDE activations caused by noise. To account for this factor in the calculation of the loss, we calculated the loss between all of the velocity estimates performed by the TDE and stimulus ground truth at corresponding timepoints (effectively, this ground truth was zero everywhere except at the moment of edge appearance see Section 5.1). Also, to evaluate how the trained TDE can reject noise and preserve the true motion signal, we employed an additional metric called fraction of true activity (FTA) as the sum of estimated velocities caused by the stimulus (true motion signal) to the total sum of estimated velocities (caused by signal and noise).
@@ 5.2.2 Training of the spiking neural networks to code with interspike interval
Apart from spike count, another intuitive method to encode signals in SNNs is relative spike timing [50, 40]. The main benefit of such signal coding is that it allows for much shorter latency [40, 50] while also enhancing coding capacity on a network level [51] However, to the best of our knowledge in the literature there is no ready solution for training neurons to have a particular ISI. We therefore decided to rectify this situation.
InputptSpatio-TemporalCorrelation FilterTDEsSpike traceISIvelocity $\sim 1/\text{ISI}$
@@
We developed a novel method to train SNNs to encode a signal with ISI using supervised learning with BPTT (Figure 10). Specifically, we were interested in the ISI between the first two spikes, as it should provide the shortest latency [27, 31]. The key is to estimate the time between two spikes in a differentiable manner. To do so we added to the pipeline on Figure 9 low-pass filtering of the TDE output spikes (Figure 10).The filtering yields an exponentially decaying spike trace $x(t)$ (3):
@@
Next, by taking filter output at the timestep directly preceding the second spike one can estimate ISI as $ISI=t(x)+1$. Then, by taking an inverse of it and scaling it by a proper factor ($\alpha$) one can convert it to the estimate of the input velocity $v_{\text{est}}=\frac{\alpha}{\text{ISI}}$ Finally, to train a network with BPTT one can calculate loss function using equation (5):
@@
    Scaling of the inverted ISI to estimate velocities. This can be done similarly to scaling of spike count: for the stimuli set with a wide dynamic range we multiplied the maximal inverted ISI (=1) with maximal stimuli velocity; for the stimuli set with a narrow dynamic range we also used bias and multiplied inverted spike count with the difference between maximal and minimal stimuli velocity.
@@ 5.2.3 Nuances of the application of TDEs to the real-world data
Real-world data was used only as a test dataset. The network was trained with synthetic moving bars to fit its dynamic range to the dynamic range of velocities in real-world data. Compared to simulation, estimation of the optical flow with TDEs in the real-world data has three additional challenges: (1) other (not only background activity) types of noise, (2) the necessity to coordinate the activity of multiple TDEs, and (3) the high dynamic range of velocities (2 decades for the rotating disk).
Apart from background activity noise, another prominent source of noise in the event-based data is jitter in event timing [18]. Event arrival is specified with microsecond precision, yet the time between actual change in intensity and event generation can vary in a range of milliseconds. Therefore, strictly speaking, each pixel of a moving edge generates an event at a different time [18]. This is especially problematic for the 3-point TDE because the prevention of its activation upon motion in an orthogonal direction is underpinned by simultaneous activation of facilitator and inhibitor, which is hardly possible given the event jitter. To avoid this problem, one needs to increase the co-activation of the facilitator and inhibitor. The most straightforward way to do so is to downsample the event stream in time. For instance, feeding events to the network using time bins, where events within a bin are considered to occur simultaneously. There is a trade-off between mitigation of event jitter and dynamic range of velocity coding. Here (Figures 7, 8), given the relatively low stimulus velocity, we chose to slice events into bins of 50 ms (i.e. fps=20).
@@
The specifics of our implementation are as follows. Using synthetic data, we first pre-trained TDE with spacing between compartments of 1 px to linearly encode 5 levels of velocity (from 0.1 px/timestep to 0.5 px/timestep). Now, it is important to recognize that the TDE encodes only the delay between activations of the facilitator and trigger compartments. Hence, the results of the training are indifferent towards the spatial configuration of the TDE, i.e. resultant velocity coding is not so much about px/timestep as about a fraction of the distance between facilitator and trigger covered within a timestep. Hence, by simply regulating the distance between the TDE inputs by a factor of n, one can scale the TDE range of velocity by said n without the need to re-train the TDE. Here, we used TDEs with distances between inputs of 1, 2, 3, 4, 6, and 8 pixels (Figure 12). Paired with a timestep duration of 50ms, it means that together the TDEs were able to cover velocities in the range of 2 to 80 px/s. For the spike count inference, the window over which spikes were counted was limited to 5 timesteps.
@@ 5.2.4 Performance evaluation for the real-world data
For the qualitative comparison between the optical flow inferred with the TDEs and the ground truth, we visualized the optical flow using the colormap in Figure 12. To make quantitative comparisons, we used the Average Angular Error (AAE), Average Endpoint Error (AEE), relative average endpoint error (rAEE), and Pearson correlation coefficient [42]. AAE measured the difference between the direction of motion estimated with the TDEs and the ground truth motion direction, and it was calculated with equation 6:
@@
The problem of AEE is that it is hard to get from it an intuition about the quality of velocity estimate. First of all, this error accumulates over time such that it becomes hard to compare algorithms run at different simulation frequencies. Secondly, understanding of whether coding error is acceptable does not depend on the absolute value of error, but on the ratio between error and true velocity. Therefore, we used rAEE to normalize AEE by the ground truth velocity to give a relative measure of performance quality using equation 8:
@@ 6 Acknowledgments
This work was supported by the Sony Research Award Program awarded to Prof. Guido C. H. E. de Croon. We want to thank Dr. Kirk Scheper for the fruitful discussions.

### v2

@@ 5.2.1 Training and inference with the synthetic data
The SNNs were simulated using the PyTorch framework, inspired by previous implementations [43, 4]. Depending on the application, networks included several types of elements such as TDEs, STCFs, and low-pass filters of the spike trains (i.e. spike traces). Each element was defined by its synaptic weights and time constants. The TDE trainable parameters included gain, current, and voltage time constants and synaptic weight of the gain compartment. To enhance learning of the neuronal time constants, we used the approach developed by [66] that reformulated neuronal decays as the result of multiplying gain, current, and voltage with the Sigmoid of a trainable parameter.
InputSpatio-TemporalCorrelation Filter TDEsVelocity = $\Sigma$ spikes
@@
Depending on the coding scheme (spike count or ISI), we used two slightly different training pipelines to extract motion direction and velocity (Figures 12,13). The two first steps are common to both pipelines. The input layer simply converts events to spikes. Then, these spikes are fed into the layer of TDE-3s, which effectively perform 1D convolution with kernel size=3, stride=1, and weights being facilitator, trigger, and inhibitor. This layer consists of two channels, one for each motion direction. At the next processing stage, spike count-based and ISI-based pipelines deviate. For the spike count-based velocity estimator we can straightaway find the total number of spikes emitted by each of the detectors, convert it to estimated velocity, calculate the loss, and perform a backward pass. For the ISI-based velocity estimation, the training of the network is not so straightforward and is described in section 5.2.2.
@@
In the experiments with noise (Figures 6, 7), not every increase in current is caused by a moving edge, and one of the goals of the training is to suppress the TDE activations caused by noise. To account for this factor in the calculation of the loss, we calculated the loss between all of the velocity estimates performed by the TDE and stimulus ground truth at corresponding timepoints (effectively, this ground truth was zero everywhere except at the moment of edge appearance see Section 5.1). Also, to evaluate how the trained TDE can reject noise and preserve the true motion signal, we employed an additional metric called fraction of true activity (FTA) as the sum of estimated velocities caused by the stimulus (true motion signal) to the total sum of estimated velocities (caused by signal and noise).
@@ 5.2.2 Training of the spiking neural networks to code with interspike interval
Apart from spike count, another intuitive method to encode signals in SNNs is relative spike timing [69, 49]. The main benefit of such signal coding is that it allows for much shorter latency [49, 69] while also enhancing coding capacity on a network level [70]. Training of the SNNs to have spike timing has a long history, with the first classical paper being published as early as 2002 [45].
InputSpatio-Temporal Correlation FilterTDEsSpike traceISIvelocity $\sim 1/\text{ISI}$
@@
Here we developed a novel method to train SNNs to encode a signal with ISI using supervised learning with BPTT (Figure 13). While most other methods [45, 46, 47] model the relationship between input magnitude with differentiable function than use BPTT, we measure ISI from the amplitude of low-pass filtered spike train to use BPTT with surrogate gradient. The theoretical advantage of our method is that it does not require estimation of the relationship between spike timing and input magnitude, allowing work with more conventional activation functions.
Specifically, we were interested in the ISI between the first two spikes, as it should provide the shortest latency [29, 33]. The key is to estimate the time between two spikes in a differentiable manner. To do so we added to the pipeline on Figure 12 low-pass filtering of the TDE output spikes (Figure 13).The filtering yields an exponentially decaying spike trace $x(t)$ (3):
@@
Next, by taking filter output at the timestep directly preceding the second spike one can estimate ISI as $ISI=t(x)+1$. Then, by taking an inverse of it and scaling it by a proper factor ($\alpha$) one can convert it to the estimate of the input velocity $v_{\text{est}}=\frac{\alpha}{\text{ISI}}$. Finally, to train a network with BPTT one can calculate the loss function using equation (5):
@@
    Scaling of the inverted ISI to estimate velocities. This can be done similarly to scaling of spike count: for the stimuli set with a wide dynamic range we multiplied the maximal inverted ISI (=1) with maximal stimuli velocity; for the stimuli set with a narrow dynamic range, we also used bias and multiplied inverted spike count with the difference between maximal and minimal stimuli velocity.
@@ 5.2.3 Nuances of the application of TDEs to the real-world data
Real-world data was used only as a test dataset. We pooled together ON and OFF events as it is often encountered in biological systems [15, 54, 55]. The network was trained with synthetic moving bars to fit its dynamic range to the dynamic range of velocities in real-world data. Compared to simulation, estimation of the optical flow with TDEs in the real-world data has three additional challenges: (1) other (not only background activity) types of noise, (2) the necessity to coordinate the activity of multiple TDEs, and (3) the high dynamic range of velocities (2 decades for the rotating disk).
Apart from background activity noise, another prominent source of noise in the event-based data is jitter in event timing [20]. Event arrival is specified with microsecond precision, yet the time between actual change in intensity and event generation can vary in a range of milliseconds. Therefore, strictly speaking, each pixel of a moving edge generates an event at a different time [20]. This is especially problematic for the 3-point TDE because the prevention of its activation upon motion in an orthogonal direction is underpinned by simultaneous activation of facilitator and inhibitor, which is hardly possible given the event jitter. To avoid this problem, one needs to increase the co-activation of the facilitator and inhibitor. The most straightforward way to do so is to downsample the event stream in time. For instance, feeding events to the network using time bins, where events within a bin are considered to occur simultaneously. There is a trade-off between mitigation of event jitter and dynamic range of velocity coding. Here (Figures 10, 11), given the relatively low stimulus velocity, we chose to slice events into bins of 50 ms (i.e. 20 frames per second).
@@
The specifics of our implementation are as follows. Using synthetic data, we first pre-trained TDE with spacing between compartments of 1 px to linearly encode 5 levels of velocity (from 0.1 px/timestep to 0.5 px/timestep). Now, it is important to recognize that the TDE encodes only the delay between activations of the facilitator and trigger compartments. Hence, the results of the training are indifferent towards the spatial configuration of the TDE, i.e. resultant velocity coding is not so much about px/timestep as about a fraction of the distance between facilitator and trigger covered within a timestep. Hence, by simply regulating the distance between the TDE inputs by a factor of n, one can scale the TDE range of velocity by said n without the need to re-train the TDE. Here, we used TDEs with distances between inputs of 1, 2, 3, 4, 6, and 8 pixels (Figure 9). Paired with a timestep duration of 50ms, it means that together the TDEs were able to cover velocities in the range of 2 to 80 px/s. For the spike count inference, the window over which spikes were counted was limited to 5 timesteps. Before feeding into the layer of TDEs signal was filtered by 3 by 3 STCF filter [18].
@@ 5.2.4 Performance evaluation for the real-world data
For the qualitative comparison between the optical flow inferred with the TDEs and the ground truth, we visualized the optical flow using the colormap in Figure 9. To make quantitative comparisons, we used the Angular Average Error (AAE), Average Endpoint Error (AEE), relative Average Endpoint Error (rAEE), and Pearson correlation coefficient [51]. AAE measured the difference between the direction of motion estimated with the TDEs and the ground truth motion direction, and it was calculated with equation 6:
@@
The problem of AEE is that it is hard to get from it an intuition about the quality of velocity estimate. First of all, this error accumulates over time such that it becomes hard to compare algorithms run at different simulation frequencies. Secondly, understanding whether coding error is acceptable does not depend on the absolute value of error, but on the ratio between error and true velocity. Therefore, we used rAEE to normalize AEE by the ground truth velocity to give a relative measure of performance quality using equation 8:
@@ 6 Acknowledgments
This work was supported by the Sony Research Award Program awarded to Prof. Guido C. H. E. de Croon.
We want to thank Dr. Kirk Scheper for the fruitful discussions.
7 Data Availability Statement
The data and code that support the findings of this study are openly available in the GitHub repository at https://github.com/yedutenko/TDE-3.